# GitHub is an effective platform for collaborative and reproducible laboratory research

### v1

@@ Abstract
Laboratory research is a complex, collaborative process that involves several stages, including hypothesis formulation, experimental design, data generation and analysis, and manuscript writing. Although reproducibility and data sharing are increasingly prioritized at the publication stage, integrating these principles at earlier stages of laboratory research has been hampered by the lack of broadly applicable solutions. Here, we propose that the workflow used in modern software development offers a robust framework for enhancing reproducibility and collaboration in laboratory research. In particular, we show that GitHub, a platform widely used for collaborative software projects, can be effectively adapted to organize and document all aspects of a research project’s lifecycle in a molecular biology laboratory. We outline a three-step approach for incorporating the GitHub ecosystem into laboratory research workflows: 1. designing and organizing experiments using issues and project boards, 2. documenting experiments and data analyses with a version control system, and 3. ensuring reproducible software environments for data analyses and writing tasks with containerized packages. The versatility, scalability, and affordability of this approach make it suitable for various scenarios, ranging from small research groups to large, cross-institutional collaborations. Adopting this framework from a project’s outset can increase the efficiency and fidelity of knowledge transfer within and across research laboratories. An example GitHub repository based on the above approach is available at https://github.com/rasilab/github_demo.
@@ Introduction
Scientific progress is contingent on the ability to reproduce and build upon previous findings. To promote reproducibility of published studies, journals and funding agencies increasingly require researchers to make their data and analysis methods available in public repositories13. In parallel, data-intensive fields such as machine learning, computational biology, and ecology have developed workflows and tools that facilitate computational reproducibility47. Reproducibility is also a critical element of collaborative research, where multiple researchers need to share and build upon each other’s work. Indeed, many reproducibility standards and tools have been developed in the context of large-scale, cross-institutional collaborations811.
@@
Compared to computational research, reproducibility and collaboration are less emphasized during the course of laboratory research, especially in the context of small research groups. This is perhaps because projects within a laboratory are often executed by individuals, and are rarely framed as collaborative efforts. However, viewed from a broader perspective, collaboration is an integral and ubiquitous feature of even small laboratory research groups. Group members discuss ideas, exchange protocols, and share reagents and data on a daily basis. Trainees collaborate with group leaders in hypothesis formulation, data interpretation, literature review, and manuscript writing. Most importantly, individuals “collaborate” with their future selves by analyzing and building on their own results, and also with future lab members who extend their work after they have left the lab. Many publicized instances of post-publication irreproducibility arise from the inability of either the same scientist at a later time or a different scientist within the same laboratory to replicate previous results12,13. Thus, workflows that improve scientific documentation and collaboration within laboratory groups from the outset of a project can enhance reproducibility of published studies across the wider research community.
@@
While much effort has been devoted to improving reproducibility during data analysis and post-publication stages of research, these do not address earlier critical stages of laboratory research. Even in a ‘wet’ lab where experiments are physically performed, researchers spend significant time and resources on literature review, hypothesis formulation, experimental design, in addition to the generation, visualization and interpretation of data. Nevertheless, few tools and workflows exist to document and share these activities in a structured and reproducible manner. Lab notebooks are the most common media used to document laboratory research, but they are typically only used for recording methods and data. An NIH handbook on lab notebooks, for example, explicitly discourages the inclusion of speculative ideas or informal conversations14, despite the recognition that these are often the source of scientific breakthroughs. Electronic lab notebooks (ELNs), despite their popularity, are stored in proprietary formats, incur a recurrent cost, tend to become defunct over time, and have poor interoperability with each other15. Cloud-based tools like Google Docs, Dropbox, and Sharepoint allow sharing of data and documents, but do not provide a structured way to track changes over time or record project-related communication. Email and messaging tools such as Slack and Microsoft Teams facilitate informal discussion of ideas and data, but these are poorly suited for organizing data and discussion in a reproducible manner. Thus, it is increasingly common for data, experiment details, analysis code, and project communication to be fragmented across multiple tools within laboratory research groups (Figure 1).
@@
The process of software development bears several similarities to activities in laboratory research. Developing software involves understanding the state-of-the-art solutions to a problem, formulating a hypothesis for improving existing approaches, breaking down the hypothesis into smaller testable features, writing code to implement features, analyzing the outcome of each feature development, and troubleshooting and fixing errors when necessary. Software development occurs in groups of all sizes ranging from a single developer to thousands of contributors, often distributed across different timezones, working on a common project. The necessity to document and share all stages of software development between contributors has led to the emergence of mature tools and workflows that facilitate reproducibility and collaboration. These include concepts such as issue tracking16, version control17, and containerization18 that have become integral to most software development projects.
Many of the common workflows associated with software development are implemented in GitHub, a popular cloud-based platform used by over 100 million developers worldwide and 90% of Fortune 100 companies19,20. GitHub repositories are highly scalable, with both small projects built by single developers and large projects with over 2,000 contributors21. In the scientific community, GitHub is used to share data analysis workflows after publication22, develop and share computational tools23, perform individual record keeping10,24,25, and conduct open science projects and collaborations11. However, how the standard workflows and rich features of GitHub (Figure 1) can be adapted to improve reproducibility and collaboration within a traditional laboratory research group has not been explored.
@@
The GitHub ‘issues’ feature provides an intuitive and flexible interface to organize and collaborate on all aspects of a laboratory experiment. In software projects, issues were originally used to track bugs or problems (hence the name ‘issue’), but their utility has expanded to new feature proposals, maintenance tasks, and general discussion topics16. Similarly, in laboratory research, we use issues not just for troubleshooting, but for organizing and discussing all aspects of research, from hypothesis formulation and experiment design to data interpretation and manuscript writing. Each issue is limited to a single topic, which focuses the ensuing discussion and resolution. In GitHub, issues are given a unique number and a URL (https://github.com/GROUP_NAME/PROJECT_NAME/issues/ISSUE_NUMBER) that provide a centralized location to track all work and discussion related to that issue. Each issue has a description field and optional commenting fields, which can be used to write, attach files, and paste images.
@@
Git is a version control system that records the history of file additions and modifications in a folder, and is used by over 90% of programmers worldwide to track changes to their code20. Git allows multiple copies of the folder to be asynchronously edited across computers, and GitHub repositories are essentially remote copies of a local folder tracked using Git. Anyone with access to a GitHub repository can download a local copy of the folder (‘clone’ in Git terminology) , add or edit file in the folder, choose which files they want to track (‘stage’), create a snapshot of the changes (‘commit’), and synchronize with the GitHub repository (‘push’). These Git features are tightly integrated into popular text editors like Visual Studio Code, which allows users to make, stage, commit, and push changes to a GitHub repository without leaving the text editor. Each commit is accompanied by a ‘commit message’ which is a short description of what changes were made since the previous commit. Importantly, the commit history of a project serves as an audit trail, recording who did what, and when.
@@
We also store data analysis materials within the GitHub project repository, which ensures that experiment logs (lab notebooks) and analysis scripts are tightly linked and easily referenced between documents. Data analysis scripts, summary tables, and visualization figures are stored in an ‘analysis’ subfolder of the parent repository, using the same hierarchical structure and naming convention as for lab notebook entries. Separate folders are created for each issue with subfolders for data, scripts, figures, tables, and sample metadata. While we store small datasets in the GitHub repository as comma- or tab-separated text files, larger datasets are stored either in private Amazon Web Services S3 folders, or in public repositories such as the Sequence Read Archive. We include short scripts to download data from their long-term storage location, which also serves as a record of the location of the data. A comma-delimited, sample metadata file is created for every dataset, following tidy principles26, to facilitate data analyses. Summary figures and tables are linked from the lab notebook page as a record of how the data was analyzed, and are linked from issue comments during data interpretation and troubleshooting discussions.
@@
The problem of replicating data analysis workflows and software environments is a common challenge in software development as well as in laboratory research. Computational researchers have long recognized this problem and have relied on tools such as the Conda package manager27, the Snakemake workflow manager28, and Docker containers to address it6. Containers are encapsulated, self-sufficient units that contain all the software needed to run an analysis, and can be shared and run on any computer that supports the container runtime environment. Public container registries, like Docker Hub or Biocontainers29, provide reproducibly-created containers, which can be used in data analysis workflows without the need to install any software. However, laboratory research groups have been slow to adopt these software reproducibility tools, which are often perceived as too complex or time-consuming to learn and use.

### v2

@@ Abstract
Laboratory research is a complex, collaborative process that involves several stages, including hypothesis formulation, experimental design, data generation and analysis, and manuscript writing. Although reproducibility and data sharing are increasingly prioritized at the publication stage, integrating these principles at earlier stages of laboratory research has been hampered by the lack of broadly applicable solutions. Here, we propose that the workflow used in modern software development offers a robust framework for enhancing reproducibility and collaboration in laboratory research. In particular, we show that GitHub, a platform widely used for collaborative software projects, can be effectively adapted to organize and document all aspects of a research project’s lifecycle in a molecular biology laboratory. We outline a three-step approach for incorporating the GitHub ecosystem into laboratory research workflows: 1. designing and organizing experiments using issues and project boards, 2. documenting experiments and data analyses with a version control system, and 3. ensuring reproducible software environments for data analyses and writing tasks with containerized packages. The versatility, scalability, and affordability of this approach make it suitable for various scenarios, ranging from small research groups to large, cross-institutional collaborations. Adopting this framework from a project’s outset can increase the efficiency and fidelity of knowledge transfer within and across research laboratories. An example GitHub repository based on this approach is available at https://github.com/rasilab/github_demo and a template repository that can be copied is available at https://github.com/rasilab/github_template.
@@ Introduction
Scientific progress is contingent on the ability to reproduce and build upon previous findings. To promote reproducibility of published studies, journals and funding agencies increasingly require researchers to make their data and analysis methods available in public repositories13. In parallel, data-intensive fields such as machine learning, computational biology, and ecology have developed workflows and tools that facilitate computational reproducibility47. Reproducibility is also a critical element of collaborative research, where multiple researchers need to share and build upon each other’s work. Indeed, many reproducibility standards and tools have been developed in the context of large-scale, cross-institutional collaborations812.
@@
Compared to computational research, reproducibility and collaboration are less emphasized during the course of laboratory research, especially in the context of small research groups. This is perhaps because projects within a laboratory are often executed by individuals, and are rarely framed as collaborative efforts. However, viewed from a broader perspective, collaboration is an integral and ubiquitous feature of even small laboratory research groups. Group members discuss ideas, exchange protocols, and share reagents and data on a daily basis. Trainees collaborate with group leaders in hypothesis formulation, data interpretation, literature review, and manuscript writing. Most importantly, individuals “collaborate” with their future selves by analyzing and building on their own results, and also with future lab members who extend their work after they have left the lab. Many publicized instances of post-publication irreproducibility arise from the inability of either the same scientist at a later time or a different scientist within the same laboratory to replicate previous results13,14. Thus, workflows that improve scientific documentation and collaboration within laboratory groups from the outset of a project can enhance reproducibility of published studies across the wider research community.
@@
While much effort has been devoted to improving reproducibility during data analysis and post-publication stages of research, these do not address earlier critical stages of laboratory research. Even in a ‘wet’ lab where experiments are physically performed, researchers spend significant time and resources on literature review, hypothesis formulation, experimental design, in addition to the generation, visualization and interpretation of data. Nevertheless, few tools and workflows exist to document and share these activities in a structured and reproducible manner. Lab notebooks are the most common media used to document laboratory research, but they are typically only used for recording methods and data. An NIH handbook on lab notebooks, for example, explicitly discourages the inclusion of speculative ideas or informal conversations15, despite the recognition that these are often the source of scientific breakthroughs. Electronic lab notebooks (ELNs), despite their popularity, are stored in proprietary formats, incur a recurrent cost, tend to become defunct over time, and have poor interoperability with each other16. Cloud-based tools like Google Docs, Dropbox, and Sharepoint allow sharing of data and documents, but do not provide a structured way to track changes over time or record project-related communication. Email and messaging tools such as Slack and Microsoft Teams facilitate informal discussion of ideas and data, but these are poorly suited for organizing data and discussion in a reproducible manner. Thus, it is increasingly common for data, experiment details, analysis code, and project communication to be fragmented across multiple tools within laboratory research groups.
@@
The process of software development bears several similarities to activities in laboratory research (Figure 1). Developing software involves understanding the state-of-the-art solutions to a problem, formulating a hypothesis for improving existing approaches, breaking down the hypothesis into smaller testable features, writing code to implement features, analyzing the outcome of each feature development, and troubleshooting and fixing errors when necessary. Software development occurs in groups of all sizes ranging from a single developer to thousands of contributors, often distributed across different timezones, working on a common project. The necessity to document and share all stages of software development between contributors has led to the emergence of mature tools and workflows that facilitate reproducibility and collaboration. These include concepts such as issue tracking17, version control18, and containerization19 that have become integral to most software development projects.
Many of the common workflows associated with software development are implemented in GitHub, a popular cloud-based platform used by over 100 million developers worldwide and 90% of Fortune 100 companies20,21. GitHub repositories are highly scalable, with both small projects built by single developers and large projects with over 2,000 contributors22. In the scientific community, GitHub is used to share data analysis workflows after publication23,24, develop and share computational tools25, perform individual record keeping10,26,27, conduct open science projects11,12, and collaboratively write manuscripts10,28. However, how the standard workflows and rich features of GitHub (Figure 1) can be adapted to improve reproducibility and collaboration within a traditional laboratory research group has not been explored.
@@
The GitHub ‘issues’ feature provides an intuitive and flexible interface to organize and collaborate on all aspects of a laboratory experiment. In software projects, issues were originally used to track bugs or problems (hence the name ‘issue’), but their utility has expanded to new feature proposals, maintenance tasks, and general discussion topics17. Similarly, in laboratory research, we use issues not just for troubleshooting, but for organizing and discussing all aspects of research, from hypothesis formulation and experiment design to data interpretation and manuscript writing. Each issue is limited to a single topic, which focuses the ensuing discussion and resolution. In GitHub, issues are given a unique number and a URL (https://github.com/GROUP_NAME/PROJECT_NAME/issues/ISSUE_NUMBER) that provide a centralized location to track all work and discussion related to that issue. Each issue has a description field and optional commenting fields, which can be used to write, attach files, and paste images.
@@
Git is a version control system that records the history of file additions and modifications in a folder, and is used by over 90% of programmers worldwide to track changes to their code21. Git allows multiple copies of the folder to be asynchronously edited across computers, and GitHub repositories are essentially remote copies of a local folder tracked using Git. Anyone with access to a GitHub repository can download a local copy of the folder (‘clone’ in Git terminology) , add or edit file in the folder, choose which files they want to track (‘stage’), create a snapshot of the changes (‘commit’), and synchronize with the GitHub repository (‘push’). These Git features are tightly integrated into popular text editors like Visual Studio Code, which allows users to make, stage, commit, and push changes to a GitHub repository without leaving the text editor. Each commit is accompanied by a ‘commit message’ which is a short description of what changes were made since the previous commit. Importantly, the commit history of a project serves as an audit trail, recording who did what, and when.
@@
We also store data analysis materials within the GitHub project repository, which ensures that experiment logs (lab notebooks) and analysis scripts are tightly linked and easily referenced between documents. Data analysis scripts, summary tables, and visualization figures are stored in an ‘analysis’ subfolder of the parent repository, using the same hierarchical structure and naming convention as for lab notebook entries. Separate folders are created for each issue with subfolders for data, scripts, figures, tables, and sample metadata. While we store small datasets in the GitHub repository as comma- or tab-separated text files, larger datasets are stored either in private Amazon Web Services S3 folders, or in public repositories such as the Sequence Read Archive. We include short scripts to download data from their long-term storage location, which also serves as a record of the location of the data. A comma-delimited, sample metadata file is created for every dataset, following tidy principles29, to facilitate data analyses. Summary figures and tables are linked from the lab notebook page as a record of how the data was analyzed, and are linked from issue comments during data interpretation and troubleshooting discussions.
@@
The problem of replicating data analysis workflows and software environments is a common challenge in software development as well as in laboratory research. Computational researchers have long recognized this problem and have relied on tools such as the Conda package manager30, the Snakemake workflow manager31, and Docker containers to address it6. Containers are encapsulated, self-sufficient units that contain all the software needed to run an analysis, and can be shared and run on any computer that supports the container runtime environment. Public container registries, like Docker Hub or Biocontainers32, provide reproducibly-created containers, which can be used in data analysis workflows without the need to install any software. However, laboratory research groups have been slow to adopt these software reproducibility tools, which are often perceived as too complex or time-consuming to learn and use.